OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i

OpenStudy (anonymous):

Here is what I have so far... (x-5),(x-3),(x-(-1+3i)),(x-(-1-3i)) x^2-8x+15 (x-(-1+3i))(x-(-1-3i))

OpenStudy (anonymous):

@dumbcow Could you help me work this out please?

OpenStudy (anonymous):

@undeadknight26 Could you help me out?

OpenStudy (anonymous):

@dumbcow I could use some guidance to answer this question, I would really appreciate it.

OpenStudy (anonymous):

Could you walk me through this problem please @Mimi_x3

OpenStudy (dumbcow):

yes you have the right idea, it can get tedious to distribute all the "i's " here is my suggestion for these problems: set x equal to the given zero, then rearrange equation so it equals 0, then you will get each factor x = 5 -----> x -5 = 0 x = -3 -----> x +3 = 0 x = -1 + 3i x+1 = 3i (square both sides to get rid of i, i^2 = -1) (x+1)^2 = -9 -----> x^2 +2x +10 = 0

OpenStudy (anonymous):

So would I do the same thing with x = -1 - 3i? @dumbcow